OpenStudy (anonymous):

lost! question is find the derivative of y with respect to x of y=(x^5/10)lnx-(x^5/50). i have the answer but i can't seem to figure out how they got it. please help!!

OpenStudy (anonymous):

dy/dx = 5x^4/10 * lnx + (x^5/10)*(1/x) - 5x^4/50 = (x^4)/2*lnx + (x^4)/10 - (x^4)/10 = (x^4)/2 * lnx